# Search-based DNN Testing and Retraining with GAN-enhanced Simulations

### v1

@@ IV Empirical Evaluation
IV-A Subjects of the Study
As subject of our study, we consider the DeeplabV3 DNN from the Gluoncv library [22], which is a state-of-the-art semantic segmentation DNN that classifies each pixel image into a predefined class. Semantic segmentation DNNs play a critical role in the development and operation of autonomous driving systems, for example, since they can be used in identifying potential obstacles such as pedestrians, other vehicles, and unexpected objects on the road; therefore, their accuracy with real (and realistic) images is key for the development of reliable safety-critical systems, a key objective for the software engineering community.
DeeplabV3 was initially trained by its developers on the Cityscapes dataset [43], which contains a variety of urban street scenes from different European cities. The dataset consists of a variety of images captured at a resolution of 1024 by 2048 pixels in the streets of 50 different European cities, primarily in Germany. Each real-world image comes with a segmentation mask representing its ground truth. The DeeplabV3 training set consists of 2975 images; however, its test set comes without ground truth and, therefore, for testing, we rely on the Cityscapes validation set (500 images). The average metric (see Section III) is particularly high, on average 80.1% (min=65%, first-quartile=73.5%, third-quartile=87.2%, max=92.4%).
To achieve the creation of images that closely mimic the appearance of real-world scenes, we utilize the Pix2PixHD model, after training it on the training set of DeeplabV3, as anticipated in Section III-A.
IV-B Assessed techniques
In our experiments, we assess a DESIGNATE prototype, that relies on AirSim 1.8.1 [36], and the original pix2pixHD GAN [37]. For DESIGNATE, we extended the NSGA-II algorithm provided by the Pymoo library [7].
To address our RQs, we compare DESIGNATE with a random baseline, two variants of DESIGNATE (for the purpose of performing ablation studies), and TACTIC [32].
Our random baseline relies only on the GAN-based input generation component of DESIGNATE; specifically, it generates realistic images from simulator segmentation maps generated using randomly selected values for the simulator parameters (e.g., position of the car in Airsim). For brevity, we refer to the images from our random baseline as random realistic images. We generate as many random images as DESIGNATE (i.e., ).
@@ IV-B Assessed techniques
TACTIC is a state-of-the-art DNN testing approach that leverages a MUNIT GAN trained to alter input images by changing their weather conditions; further, it integrates a search-based approach which alters the latent space vector in such a way that the generated images lead to the worst accuracy for the DNN under test. To apply TACTIC, we randomly select 220 images from the Cityscapes validation set and apply the five weather conditions supported by TACTIC (night, rain, snow night, snow daytime, and sunny conditions), each leading to a distinct new test image. This process generates a total of 1100 images.
@@ IV-C1 Design and measurements
For each assessed technique, we compute the obtained by the DeepLabV3 model with the generated images. Since we aim to discover critical DNN limitations, the most effective technique is the one leading to the lowest median, minimum, and first quartile.
However, since we also wish to obtain a set of diverse unsafe situations, we also assess diversity across the generated images by relying on the two distance metrics employed by DESIGNATE: feature-based distance and pixel-based distance. Specifically, we compute the distance between every possible pair of images generated by the selected techniques and compare their distribution. The best approach should lead to the highest diversity value.
@@ IV-C2 Results
Accuracy. Table II provides information about the accuracy of the subject DNN when tested by the selected approaches. Column images provides the average number of test images selected over the ten executions; for DESIGNATE and its variants we report the images in the archive, for Random and TACTIC we report all the generated images (for Random, we excluded, on average, 15 images that do not satisfy ‘’). The other columns provide descriptive statistics of the metric. Table V compares each pair of techniques providing p-values and .
DESIGNATEsingle is the approach leading to the largest number of images with the lowest ; indeed, it shows the lowest average and median values. However, although their difference is always significant (p-value , Table V), based on , DESIGNATEsingle, DESIGNATE, and DESIGNATEpix perform similarly.
Except for DESIGNATENoGAN, all the DESIGNATE variants perform significantly better (lower ) than Random and TACTIC (Table V), which indicates that combining simulators driven by search with GAN is necessary for effective testing. For example, DESIGNATE leads to a median of compared to and for Random and TACTIC, respectively. Further, the first quartile of all the GAN-based DESIGNATE variants is zero, indicating that 25% of the generated images do not correctly identify a single pixel belonging to cars. For Random, less than 25% of the images lead to (first quartile is ), while TACTIC always lead to a high (), with a median being similar to the one obtained with Random (i.e., 0.82 vs. 0.85). Also, Table V shows that the difference between DESIGNATE and both TACTIC and Random is practically significant (non-equivalent populations according to ). DESIGNATENoGAN, instead, leads to median results () that are in-between those of TACTIC and Random, with indicating that DESIGNATENoGAN and TACTIC perform similarly, while Random performs slightly better than DESIGNATENoGAN (). These results suggest that relying on simulator images only is not sufficient to drive the search towards cases where the DNN performs worse.
@@
The random baseline appears to be the least effective, with the highest median of 85%, which shows that the DNN under test performs generally well in the sampled input space and a search-based approach is needed to comprehensively uncover its limitations.
@@
DESIGNATEsingle | DESIGNATE | DESIGNATEpix | DESIGNATENoGAN | Random | TACTIC | |||||||
p-value | p-value | p-value | p-value | p-value | p-value | |||||||
DESIGNATEsingle | 0.45 | 5.00E-28 | 0.47 | 8.00E-13 | 0.28 | 0.00E+00 | 0.34 | 3.00E-298 | 0.35 | 1.00E-64 | ||
DESIGNATE | 0.55 | 5.00E-28 | 0.52 | 1.00E-03 | 0.33 | 0.00E+00 | 0.39 | 1.14E-139 | 0.41 | 9.00E-21 | ||
DESIGNATEpix | 0.53 | 8.00E-13 | 0.48 | 2.00E-03 | 0.31 | 0.00E+00 | 0.38 | 8.00E-161 | 0.41 | 1.00E-22 | ||
DESIGNATENoGAN | 0.72 | 0.00E+00 | 0.67 | 0.00E+00 | 0.69 | 0.00E+00 | 0.57 | 0.00E+00 | 0.55 | 0.00E+00 | ||
Random | 0.66 | 3.00E-298 | 0.61 | 1.14E-135 | 0.62 | 8.00E-161 | 0.43 | 0.00E+00 | 0.54 | 5.10E-05 | ||
TACTIC | 0.64 | 1.00E-64 | 0.58 | 9.00E-21 | 0.59 | 1.00E-22 | 0.45 | 0.00E+00 | 0.46 | 5.10E-05 |
Note: tangible (based on ) differences in bold. the best approach is the one leading to the lowest .
@@
Feature-based distance | ||||||||||||
@@
DESIGNATEsingle | DESIGNATE | DESIGNATEpix | DESIGNATENoGAN | Random | TACTIC | |||||||
p-value | p-value | p-value | p-value | p-value | p-value | |||||||
DESIGNATEsingle | 0.33 | 0.00E+00 | 0.51 | 0.00E+00 | 0.55 | 0.00E+00 | 0.78 | 0.00E+00 | 0.88 | 0.00E+00 | ||
DESIGNATE | 0.67 | 0.00E+00 | 0.67 | 0.00E+00 | 0.72 | 0.00E+00 | 0.86 | 0.00E+00 | 0.93 | 0.00E+00 | ||
DESIGNATEpix | 0.49 | 0.00E+00 | 0.33 | 0.00E+00 | 0.53 | 0.00E+00 | 0.75 | 0.00E+00 | 0.85 | 0.00E+00 | ||
DESIGNATENoGAN | 0.45 | 0.00E+00 | 0.28 | 0.00E+00 | 0.47 | 0.00E+00 | 0.77 | 0.00E+00 | 0.84 | 0.00E+00 | ||
Random | 0.22 | 0.00E+00 | 0.14 | 0.00E+00 | 0.25 | 0.00E+00 | 0.23 | 0.00E+00 | 0.74 | 9.00E-292 | ||
TACTIC | 0.12 | 0.00E+00 | 0.07 | 0.00E+00 | 0.15 | 0.00E+00 | 0.16 | 0.00E+00 | 0,26 | 9.00E-292 | ||
Pixel-distance | ||||||||||||
@@
DESIGNATEsingle | DESIGNATE | DESIGNATEpix | DESIGNATENoGAN | Random | TACTIC | |||||||
p-value | p-value | p-value | p-value | p-value | p-value | |||||||
DESIGNATEsingle | 0.39 | 0.00E+00 | 0.48 | 0.00E+00 | 0.95 | 0.00E+00 | 0.99 | 0.00E+00 | 0.43 | 0.00E+00 | ||
DESIGNATE | 0.61 | 0.00E+00 | 0.58 | 0.00E+00 | 0.96 | 0.00E+00 | 0.99 | 0.00E+00 | 0.58 | 1.00E-21 | ||
DESIGNATEpix | 0.52 | 0.00E+00 | 0.42 | 0.00E+00 | 0.94 | 0.00E+00 | 0.99 | 0.00E+00 | 0.46 | 3.00E-07 | ||
DESIGNATENoGAN | 0.05 | 0.00E+00 | 0.04 | 0.00E+00 | 0.06 | 0.00E+00 | 1.00 | 0.00E+00 | 0.01 | 0.00E+00 | ||
Random | 0.01 | 0.00E+00 | 0.01 | 0.00E+00 | 0.01 | 0.00E+00 | 0.00 | 0.00E+00 | 0.00 | 0.00E+00 | ||
TACTIC | 0.57 | 0.00E+00 | 0.42 | 1.00E-21 | 0.54 | 3.00E-07 | 0.99 | 0.00E+00 | 1.00 | 0.00E+00 |
@@
Note: tangible (based on ) differences in bold, the best approach is the one leading to the highest .
@@
Diversity. Table IV provides descriptive statistics for the diversity metrics collected for the selected techniques. Table V provides p-values and resulting from the comparisons of diversity across techniques.

### v2

@@ IV-B Assessed techniques
TACTIC is a state-of-the-art DNN testing approach that leverages a MUNIT GAN trained to alter input images by changing their weather conditions (night, sunshine, rain, snow); further, it integrates a search-based approach which alters the latent space vector in such a way that the generated images lead to the worst accuracy for the DNN under test. To apply TACTIC, we randomly select 220 images from the Cityscapes validation set and apply the five weather conditions supported by TACTIC (night, rain, snow night, snow daytime, and sunny conditions), each leading to a distinct new test image. This process generates a total of 1100 images. ††margin: R3.6 Since TACTIC was originally trained to generate driving conditions, we consider it for the Urban environment only; indeed, it makes little sense to add snow and rain effects on Mars images, as well as, changing illumination to make it similar to Earth.
DeepJanus is multi-objective search-based approach that searches for the boundary of the input space that is properly processed by the DNN; specifically, it looks for pair of images that are similar but leads to different DNN results, one of which is erroneous. ††margin: R2.1 We integrate DeepJanus with our simulators with the same configurations as in the original paper. Furher, since DeepJanus exclusively produces simulated images, we explore the impact of using a GAN to generate more realistic images (hereafter, DeepJanusGAN). To achieve this, we incorporated the Pix2pixHD model, as utilized in DESIGNATE, into the DeepJanus workflow to transform simulated images into realistic ones; precisely, DeepJanusGAN, during search, still generates simulator images and ground truth as in DeepJanus, but, instead of providing a simulator image to the DNN under test, it first generates a realistic image using Pix2pixHD and then feeds it to the DNN under test.
@@ IV-C1 Design and measurements
††margin: R3.6For each assessed technique, we compute the , for the urban case study, and for the Martian case study (see Section III-B1 for definitions), obtained by the DeepLabV3 model with the generated images. Since we aim to discover critical DNN limitations, the most effective technique is the one leading to the lowest median, minimum, and first quartile.
However, since we also wish to obtain a set of diverse unsafe situations, we also assess diversity across the generated images by relying on the two distance metrics employed by DESIGNATE: feature-based distance and pixel-based distance. Specifically, we compute the distance between every possible pair of images generated by the selected techniques and compare their distributions. An approach with higher diversity values is preferable to identify the different situations under which segmentations are erroneous.
@@ IV-C2 Results
Accuracy. ††margin: R3.6 Table II and IV provide, for each of the selected approaches, details on the performance of the subject DNNs for the urban and the Martian environment, respectively. Column images provides the average number of test images selected over the ten executions; for DESIGNATE and its variants we report the images in the archive, for Random and TACTIC we report all the generated images, excluding, for the former, an average of 15 and 169 images, for the urban and Martin environments, respectively, that do not satisfy ‘’. The other columns provide descriptive statistics for and ††margin: R3.6 . Table VII and Table V report differences between each pair of techniques by providing p-values and for the urban and martian environments, respectively. We discuss the results observed with the urban environment first.
DESIGNATEsingle is the approach leading to the largest number of images with low values; indeed, it shows the lowest average ($0.48$) and median ($0.67$) values. However, although their difference is always statistically significant (p-value , Table VII), based on , DESIGNATEsingle, DESIGNATE, DESIGNATEpix, and DeepJanusGAN perform similarly and do not lead to practically significant differences.
Except for DESIGNATENoGAN, all the DESIGNATE variants perform significantly better (lower ) than Random, and TACTIC (Table VII), which indicates that combining simulators driven by search with GAN is necessary for effective testing. For example, DESIGNATE leads to a median of compared to and for Random and TACTIC, respectively. Further, the first quartile of all the GAN-based DESIGNATE variants is zero, indicating that 25% of the generated images do not correctly identify a single pixel belonging to cars. For Random, less than 25% of the images lead to (first quartile is ), while TACTIC tends to lead to a higher (), with never being equal to 0. Also, all the DESIGNATE variants fare better than DeepJanusGAN, thus showing that an appropriate search strategy is necessary to identify test inputs leading to failures. As for DeepJanus, its is similar in terms of median to that of DESIGNATEsingle and worse in terms of average.
Also, Table III shows that the difference between DESIGNATE and both TACTIC and Random is practically significant, according to . DESIGNATENoGAN, instead, leads to median results () that are in-between those of TACTIC and Random, with indicating that DESIGNATENoGAN and TACTIC perform similarly, while Random performs slightly better than DESIGNATENoGAN (). These results suggest that relying on simulator images only is not sufficient to drive the search towards cases where the DNN performs worse.
Concerning TACTIC, we further checked how the different weather conditions (rain, snow day, sunny, and night) perform. Figure 5 illustrates an example of a generated image for each weather condition, alongside the median of the generated images for the respective conditions. Our observations indicate that three conditions — Sunny (median = 80%, min = 50%, max = 97%, average = 78%), Night (median = 90%, min = 50%, max = 94%, average = 80%), and Snow-Night (median = 80%, min = 50%, max = 91%, average = 80%) — yield high median values, with the generated images appearing realistic. Conversely, the Rainy (median = 10%, min = 1%, max = 32%, average = 10%) and Snow-Day (median = 10%, min = 2%, max = 36%, average = 10%) conditions result in lower median values, with the generated images being unrealistic, with no observable cars, and therefore not matching the ground truth. Such images, as illustrated in Figure 5, are not useful for testing purposes.
@@
The random baseline appears to be less effective, with a high median of 85%, which shows that the DNN under test performs generally well in the sampled input space and a search-based approach is needed to comprehensively uncover its limitations.
DeepJanus yields slightly better accuracy compared to DESIGNATE and DESIGNATEpix with suggesting that DeepJanus, DESIGNATE, and DESIGNATEpix perform similarly. However, DESIGNATEsingle outperforms DeepJanus, which suggests that the single-objective search used by DESIGNATEsingle leads to better results than the one used by DeepJanus. Further, after visual inspection, we realized that most the failures triggered by DeepJanus are due to borderline cases with cars appearing far away and occupying few pixels in the image. Though such sceneries can be generated by simulators, they are likely to be irrelevant in practice as the position of cars that are far away are unlikely to affect the behaviour of the ego vehicle and thus be relevant for testing. More precisely, because DeepJanus looks for frontier behaviour, with pairs of images that are very similar but where one of the two leads to failures, it tends to generate images where cars are far from the ego vehicle.
The higher accuracy of DeepJanusGAN, which performs worse and significantly different from DESIGNATE, is likely due to the fact that, when applying GANs, far away cars disappear from the picture. Indeed, the training set does not contain cars occupying only few pixels, because such far-away cars are irrelevant for a segmentation DNN in an automotive system and thus not important for testing purposes.
For the Martian environment, DESIGNATE achieved a median of 0.49, outperforming DESIGNATEsingle (0.54), DESIGNATEpix (0.52), Random (0.64), and DeepJanusGAN (0.60). Similar to the urban environment, the values in Table V indicate that DESIGNATEsingle, DESIGNATE, and DESIGNATEpix perform similarly, although their differences are statistically significant ().
The key difference between the urban and Martian environments is that the lowest accuracy values are observed with DESIGNATENoGAN and DeepJanus: 0.32 and 0.48 in the latter, as opposed to DESIGNATEsingle in the former. We attribute this result to DESIGNATENoGAN and DeepJanus relying on simulated images that are often unrealistic, thereby increasing the likelihood of mispredictions. This can be explained by the low fidelity of the Mars simulator when compared to AirSim, as depicted in Fig. 3. To conclude, as reported for the urban environment, the use of GAN prevents the identification of failures that will not occur in real usage settings. In other words, when not using a GAN and relying on a low-fidelity simulator, many unrealistic and therefore irrelevant test inputs are generated.
@@
DESIGNATEsingle | DESIGNATE | DESIGNATEpix | DESIGNATENoGAN | Random | TACTIC | DeepJanus | DeepJanusGAN | |||||||||
p-value | p-value | p-value | p-value | p-value | p-value | p-value | p-value | |||||||||
DESIGNATEsingle | 0.45 | 5.00E-28 | 0.47 | 8.00E-13 | 0.28 | 0.00E+00 | 0.34 | 3.00E-298 | 0.35 | 1.00E-64 | 0.47 | 2.19E-09 | 0.33 | 7.53E-301 | ||
DESIGNATE | 0.55 | 5.00E-28 | 0.52 | 1.00E-03 | 0.33 | 0.00E+00 | 0.39 | 1.14E-139 | 0.41 | 9.00E-21 | 0.53 | 9.24E-09 | 0.38 | 1.36E-131 | ||
DESIGNATEpix | 0.53 | 8.00E-13 | 0.48 | 2.00E-03 | 0.31 | 0.00E+00 | 0.38 | 8.00E-161 | 0.41 | 1.00E-22 | 0.52 | 4.69E-03 | 0.38 | 2.41E-150 | ||
DESIGNATENoGAN | 0.72 | 0.00E+00 | 0.67 | 0.00E+00 | 0.69 | 0.00E+00 | 0.57 | 0.00E+00 | 0.55 | 0.00E+00 | 0.73 | 1.88E-133 | 0.61 | 1.19E-33 | ||
Random | 0.66 | 3.00E-298 | 0.61 | 1.14E-135 | 0.62 | 8.00E-161 | 0.43 | 0.00E+00 | 0.54 | 5.10E-05 | 0.66 | 2.23E-210 | 0.50 | 6.95E-01 | ||
TACTIC | 0.64 | 1.00E-64 | 0.58 | 9.00E-21 | 0.59 | 1.00E-22 | 0.45 | 0.00E+00 | 0.46 | 5.10E-05 | 0.73 | 3.40E-55 | 0.45 | 2.15E-04 | ||
DeepJanus | 0.53 | 2.19E-09 | 0.47 | 9.24E-09 | 0.48 | 4.69E-03 | 0.27 | 1.88E-133 | 0.34 | 2.23E-210 | 0.27 | 3.40E-55 | 0.29 | 2.18E-297 | ||
DeepJanusGAN | 0.67 | 7.53E-301 | 0.62 | 1.36E-131 | 0.62 | 2.41E-150 | 0.39 | 1.19E-33 | 0.50 | 6.95E-01 | 0.55 | 2.15E-04 | 0.71 | 2.18E-297 |
Note: tangible (based on ) differences in bold; for each pair of approaches, the best approach is the one with and its name on the row.
#images | ||||||||
DESIGNATEsingle | 802 | 0.01 | 0.96 | 0.54 | 0.09 | 0.35 | 0.73 | 0.52 |
DESIGNATE | 764 | 0.01 | 0.96 | 0.49 | 0.09 | 0.33 | 0.69 | 0.50 |
DESIGNATEpix | 718 | 0.01 | 0.95 | 0.52 | 0.20 | 0.39 | 0.70 | 0.53 |
DESIGNATENoGAN | 1151 | 0.01 | 0.97 | 0.32 | 0.03 | 0.12 | 0.70 | 0.39 |
Random | 1031 | 0.01 | 0.96 | 0.64 | 0.24 | 0.44 | 0.83 | 0.62 |
DeepJanus | 986 | 0.01 | 0.93 | 0.48 | 0.03 | 0.20 | 0.67 | 0.44 |
DeepJanusGAN | 954 | 0.01 | 0.98 | 0.60 | 0.20 | 0.42 | 0.85 | 0.61 |
Note: best (i.e., lowest) results (per column) in bold.
DESIGNATEsingle | DESIGNATE | DESIGNATEpix | DESIGNATENoGAN | Random | DeepJanus | DeepJanusGAN | ||||||||
p-value | p-value | p-value | p-value | p-value | p-value | p-value | ||||||||
DESIGNATEsingle | 0.53 | 1.10E-11 | 0.50 | 4.41E-01 | 0.63 | 4.55E-197 | 0.38 | 3.11E-153 | 0.59 | 1.80E-81 | 0.40 | 1.45E-116 | ||
DESIGNATE | 0.47 | 1.10E-11 | 0.46 | 5.22E-16 | 0.61 | 2.20E-136 | 0.36 | 4.16E-228 | 0.56 | 2.89E-34 | 0.37 | 5.03E-178 | ||
DESIGNATEpix | 0.50 | 0.440865115 | 0.54 | 5.22E-16 | 0.65 | 2.54E-239 | 0.38 | 6.00E-146 | 0.59 | 2.76E-85 | 0.40 | 1.90E-103 | ||
DESIGNATENoGAN | 0.37 | 4.55E-197 | 0.39 | 2.20E-136 | 0.35 | 2.54E-239 | 0.28 | 0.00E+00 | 0.46 | 3.18E-24 | 0.28 | 0.00E+00 | ||
Random | 0.62 | 3.11E-153 | 0.64 | 4.16E-228 | 0.62 | 6.00E-146 | 0.72 | 0.00E+00 | 0.69 | 0.00E+00 | 0.50 | 3.97E-01 | ||
DeepJanus | 0.41 | 1.80E-81 | 0.44 | 2.89E-34 | 0.41 | 2.76E-85 | 0.54 | 3.18E-24 | 0.31 | 0.00E+00 | 0.33 | 0.00E+00 | ||
DeepJanusGAN | 0.60 | 1.45E-116 | 0.63 | 5.03E-178 | 0.60 | 1.90E-103 | 0.72 | 0.00E+00 | 0.50 | 3.97E-01 | 0.67 | 0.00E+00 |
Note: tangible (based on ) differences in bold; for each pair of approaches, the best approach is the one with and its name on the row.
@@
Feature-based distance | ||||||||||||||||
@@
DESIGNATEsingle | DESIGNATE | DESIGNATEpix | DESIGNATENoGAN | Random | TACTIC | DeepJanus | DeepJanusGAN | |||||||||
p-value | p-value | p-value | p-value | p-value | p-value | p-value | p-value | |||||||||
DESIGNATEsingle | 0.33 | 0.00E+00 | 0.51 | 0.00E+00 | 0.55 | 0.00E+00 | 0.78 | 0.00E+00 | 0.87 | 0.00E+00 | 0.45 | 0.00E+00 | 0.57 | 0.00E+00 | ||
DESIGNATE | 0.67 | 0.00E+00 | 0.67 | 0.00E+00 | 0.72 | 0.00E+00 | 0.86 | 0.00E+00 | 0.92 | 0.00E+00 | 0.64 | 0.00E+00 | 0.72 | 0.00E+00 | ||
DESIGNATEpix | 0.49 | 0.00E+00 | 0.33 | 0.00E+00 | 0.53 | 0.00E+00 | 0.75 | 0.00E+00 | 0.84 | 0.00E+00 | 0.44 | 0.00E+00 | 0.56 | 0.00E+00 | ||
DESIGNATENoGAN | 0.45 | 0.00E+00 | 0.28 | 0.00E+00 | 0.47 | 0.00E+00 | 0.77 | 0.00E+00 | 0.84 | 0.00E+00 | 0.40 | 0.00E+00 | 0.54 | 0.00E+00 | ||
Random | 0.22 | 0.00E+00 | 0.14 | 0.00E+00 | 0.25 | 0.00E+00 | 0.23 | 0.00E+00 | 0.74 | 5.55E-169 | 0.14 | 0.00E+00 | 0.30 | 0.00E+00 | ||
TACTIC | 0.13 | 0.00E+00 | 0.08 | 0.00E+00 | 0.16 | 0.00E+00 | 0.16 | 0.00E+00 | 0.26 | 5.55E-169 | 0.06 | 0.00E+00 | 0.18 | 7.40E-299 | ||
DeepJanus | 0.55 | 0.00E+00 | 0.36 | 0.00E+00 | 0.56 | 0.00E+00 | 0.60 | 0.00E+00 | 0.86 | 0.00E+00 | 0.94 | 0.00E+00 | 0.63 | 0.00E+00 | ||
DeepJanusGAN | 0.43 | 0.00E+00 | 0.28 | 0.00E+00 | 0.44 | 0.00E+00 | 0.46 | 0.00E+00 | 0.70 | 0.00E+00 | 0.82 | 7.40E-299 | 0.37 | 0.00E+00 | ||
Pixel-based distance | ||||||||||||||||
@@
DESIGNATEsingle | DESIGNATE | DESIGNATEpix | DESIGNATENoGAN | Random | TACTIC | DeepJanus | DeepJanusGAN | |||||||||
p-value | p-value | p-value | p-value | p-value | p-value | p-value | p-value | |||||||||
DESIGNATEsingle | 0.39 | 0.00E+00 | 0.48 | 0.00E+00 | 0.95 | 0.00E+00 | 0.50 | 1.00E+00 | 0.44 | 3.77E-12 | 0.86 | 0.00E+00 | 0.57 | 0.00E+00 | ||
DESIGNATE | 0.61 | 0.00E+00 | 0.58 | 0.00E+00 | 0.96 | 0.00E+00 | 0.61 | 0.00E+00 | 0.59 | 2.74E-22 | 0.90 | 0.00E+00 | 0.67 | 0.00E+00 | ||
DESIGNATEpix | 0.52 | 0.00E+00 | 0.42 | 0.00E+00 | 0.94 | 0.00E+00 | 0.52 | 0.00E+00 | 0.46 | 3.76E-05 | 0.86 | 0.00E+00 | 0.59 | 0.00E+00 | ||
DESIGNATENoGAN | 0.05 | 0.00E+00 | 0.04 | 0.00E+00 | 0.06 | 0.00E+00 | 0.05 | 0.00E+00 | 0.01 | 0.00E+00 | 0.29 | 0.00E+00 | 0.08 | 0.00E+00 | ||
Random | 0.50 | 1.00E+00 | 0.39 | 0.00E+00 | 0.48 | 0.00E+00 | 0.95 | 0.00E+00 | 0.44 | 3.77E-12 | 0.86 | 0.00E+00 | 0.57 | 0.00E+00 | ||
TACTIC | 0.56 | 3.77E-12 | 0.41 | 2.74E-22 | 0.54 | 3.76E-05 | 0.99 | 0.00E+00 | 0.56 | 3.77E-12 | 0.96 | 0.00E+00 | 0.65 | 2.11E-64 | ||
DeepJanus | 0.14 | 0.00E+00 | 0.10 | 0.00E+00 | 0.14 | 0.00E+00 | 0.71 | 0.00E+00 | 0.14 | 0.00E+00 | 0.04 | 0.00E+00 | 0.19 | 0.00E+00 | ||
DeepJanusGAN | 0.43 | 0.00E+00 | 0.33 | 0.00E+00 | 0.41 | 0.00E+00 | 0.92 | 0.00E+00 | 0.43 | 0.00E+00 | 0.35 | 2.11E-64 | 0.81 | 0.00E+00 |
@@
Note: tangible (based on ) differences in bold; for each pair of approaches, the best approach is the one with and its name on the row.
Feature Distance Diversity | |||||||
DESIGNATEsingle | 0.00 | 27.17 | 10.59 | 5.67 | 8.70 | 12.35 | 10.58 |
DESIGNATE | 0.00 | 26.24 | 14.75 | 8.09 | 11.22 | 18.58 | 14.83 |
DESIGNATEpix | 0.82 | 25.43 | 11.92 | 7.79 | 9.99 | 14.38 | 12.37 |
DESIGNATENoGAN | 1.01 | 32.49 | 16.79 | 10.23 | 14.35 | 19.54 | 16.83 |
Random | 0.00 | 28.70 | 11.36 | 7.51 | 9.44 | 13.66 | 11.89 |
DeepJanus | 2.18 | 31.67 | 12.81 | 7.16 | 10.58 | 18.64 | 14.48 |
DeepJanusGAN | 0.00 | 25.97 | 9.52 | 6.13 | 7.80 | 12.25 | 10.54 |
Pixel Distance Diversity | |||||||
DESIGNATEsingle | 0.96 | 1.00 | 0.99 | 0.97 | 0.98 | 0.99 | 0.99 |
DESIGNATE | 0.95 | 1.00 | 1.00 | 0.99 | 1.00 | 1.00 | 1.00 |
DESIGNATEpix | 0.95 | 1.00 | 0.98 | 0.96 | 0.97 | 0.99 | 0.98 |
DESIGNATENoGAN | 0.95 | 1.00 | 0.98 | 0.96 | 0.97 | 0.99 | 0.98 |
Random | 0.96 | 1.00 | 0.98 | 0.97 | 0.98 | 0.99 | 0.98 |
DeepJanus | 0.95 | 0.99 | 0.96 | 0.95 | 0.96 | 0.97 | 0.96 |
DeepJanusGAN | 0.95 | 1.00 | 0.99 | 0.97 | 0.98 | 0.99 | 0.98 |
Note: best (i.e., highest or most diverse) results (per column) in bold.
Feature-based distance | ||||||||||||||
DESIGNATEsingle | DESIGNATE | DESIGNATEpix | DESIGNATENoGAN | Random | DeepJanus | DeepJanusGAN | ||||||||
p-value | p-value | p-value | p-value | p-value | p-value | p-value | ||||||||
DESIGNATEsingle | 0.43 | 0.00E+00 | 0.52 | 0.00E+00 | 0.15 | 0.00E+00 | 0.51 | 0.00E+00 | 0.34 | 0.00E+00 | 0.62 | 0.00E+00 | ||
DESIGNATE | 0.57 | 0.00E+00 | 0.59 | 0.00E+00 | 0.19 | 0.00E+00 | 0.59 | 0.00E+00 | 0.40 | 0.00E+00 | 0.68 | 0.00E+00 | ||
DESIGNATEpix | 0.48 | 0.00E+00 | 0.41 | 0.00E+00 | 0.13 | 0.00E+00 | 0.50 | 0.00E+00 | 0.33 | 0.00E+00 | 0.62 | 0.00E+00 | ||
DESIGNATENoGAN | 0.85 | 0.00E+00 | 0.81 | 0.00E+00 | 0.87 | 0.00E+00 | 0.86 | 0.00E+00 | 0.70 | 0.00E+00 | 0.88 | 0.00E+00 | ||
Random | 0.49 | 0.00E+00 | 0.41 | 0.00E+00 | 0.50 | 0.00E+00 | 0.14 | 0.00E+00 | 0.33 | 0.00E+00 | 0.62 | 0.00E+00 | ||
DeepJanus | 0.66 | 0.00E+00 | 0.60 | 0.00E+00 | 0.67 | 0.00E+00 | 0.30 | 0.00E+00 | 0.67 | 0.00E+00 | 0.74 | 0.00E+00 | ||
DeepJanusGAN | 0.38 | 0.00E+00 | 0.32 | 0.00E+00 | 0.38 | 0.00E+00 | 0.12 | 0.00E+00 | 0.38 | 0.00E+00 | 0.26 | 0.00E+00 | ||
Pixel-based distance | ||||||||||||||
DESIGNATEsingle | DESIGNATE | DESIGNATEpix | DESIGNATENoGAN | Random | DeepJanus | DeepJanusGAN | ||||||||
p-value | p-value | p-value | p-value | p-value | p-value | p-value | ||||||||
DESIGNATEsingle | 0.36 | 2.23E-211 | 0.60 | 3.56E-106 | 0.81 | 0.00E+00 | 0.55 | 1.08E-26 | 0.99 | 0.00E+00 | 0.55 | 1.32E-35 | ||
DESIGNATE | 0.64 | 2.23E-211 | 0.71 | 0.00E+00 | 0.83 | 0.00E+00 | 0.68 | 0.00E+00 | 0.93 | 0.00E+00 | 0.68 | 0.00E+00 | ||
DESIGNATEpix | 0.40 | 3.56E-106 | 0.29 | 0.00E+00 | 0.77 | 0.00E+00 | 0.44 | 4.51E-40 | 0.98 | 0.00E+00 | 0.45 | 0.00E+00 | ||
DESIGNATENoGAN | 0.19 | 0.00E+00 | 0.17 | 0.00E+00 | 0.23 | 0.00E+00 | 0.20 | 0.00E+00 | 0.88 | 0.00E+00 | 0.21 | 0.00E+00 | ||
Random | 0.45 | 1.08E-26 | 0.32 | 0.00E+00 | 0.56 | 4.51E-40 | 0.80 | 0.00E+00 | 0.99 | 0.00E+00 | 0.51 | 0.00E+00 | ||
DeepJanus | 0.01 | 0.00E+00 | 0.07 | 0.00E+00 | 0.02 | 0.00E+00 | 0.12 | 0.00E+00 | 0.01 | 0.00E+00 | 0.01 | 0.00E+00 | ||
DeepJanusGAN | 0.45 | 1.32E-35 | 0.32 | 0.00E+00 | 0.55 | 1.11E-34 | 0.79 | 0.00E+00 | 0.49 | 1.19E-01 | 0.99 | 0.00E+00 |
Note: tangible (based on ) differences in bold; for each pair of approaches, the best approach is the one with and its name on the row.
@@
Diversity. Table VI and Table VIII provide descriptive statistics for the diversity metrics collected for the selected techniques for the Urban and Martian environments, respectively. Additionally, Table VII and Table IX provide p-values and resulting from the comparisons of diversity across techniques for the Urban and Martian environments, respectively. We discuss the urban environment first.